Determine which of the following is the solution set of the linear equations below.


3x − y + z = 2

2x − z = 2


1
Expert's answer
2022-04-12T15:53:57-0400

The system has an infinite number of solutions because we have 3 variables and only 2 equations:

"\begin{cases}3x-y+z=2\\
2x-z=2
 \end{cases}
\implies
\begin{cases}
5x-y=4\\
2x-z=2
\end{cases}
\implies""\begin{cases}
5x-y=4\\
x=\frac12(2+z)
\end{cases}
\implies\begin{cases}
5x-y=4\\
x=1+0.5z
\end{cases}\implies" "\begin{cases}
5(1+0.5z)-y=4\\
x=1+0.5z
\end{cases}\implies\begin{cases}
5+2.5z-y=4\\
x=1+0.5z
\end{cases}""\implies\begin{cases}
y=1+2.5z\\
x=1+0.5z
\end{cases}\implies" "\begin{cases}
x=1+0.5z\\
y=1+2.5z\\
z\in R
\end{cases}"

"R" - set of real numbers.